X-GEF TESTS for Radiation Tests file: ~dd/text/VSG/rad_test.txt - - - DATA STATUS ----------- Currently the data sets are in: pspc_directory = 'cenx3/d11/tgs/YYMMDD' and backup (not touched!) copies in "YYMMDDb" directories. Key (at beginning of the lines): :( means this is a problem... ** means this is something To-Do. !- means "Eureka!" 001030 [2077]1-187: - H2102[B] and HB2121[F] (VSG gratings) :( as Bob mentioned the SSD HV was off so no SSD data - I am working to get PSPC calibrated: in X and Y it is OK, but a bit finicky in Energy calibration... Problem is PSPC is flipped OR detector translation sign flipped! - Analysis (No SSD): SUMMARY, SMD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT Tilt plot for HB2121p0tilt looks just like 990715 one! Diffraction efficiences are within a few % from 990715(VSG) test: Grating/ Test 1 keV 4 keV 8 keV HB2121/001030 p0: 2.42 16.43 9.15 HB2121/990715 p0: 2.38 16.09 8.81 /nfs/spectra/d1/990715 H2102 /001030 p0: 1.68 12.75 7.23 H2102 /990715 p0: 1.62 12.50 7.27 /nfs/spectra/d1/990715 001101 [2079]1-156: - HC2214[C] and HC2259[C] (damaged at alignment; Eng HESS) - was in directory named 001030b ==> renamed it 001101 :( Zero-length or missing PSPC items 157-160, SMD 156-160; No SSD items. Put zero-length files in '/NA_files/ - Analysis: SUMMARY, SMD, PSPC_CAL Make PSPC_ANAL, etc, work even if some data files are missing... If data are available on tape reload the tape data into --> 001101t <-- directory which exists and is empty now. - Completed analysis with the few missing files. 001113 [2083]1-157,167-187: - HA2021[A] and HB2104[B] (Synch. tested, Acoustic 96) - some SSD files were missing from this directory and I found them (?) in /nfs/caliber/dd/scratch and in 001114. :( some of the files have zero-length: the last three PSPC files and the first 6 SSD files. - This may just be a tape-reading thing and be recoverable? If so, put data-from-tape in 001113b or make an 001113c ... - I saw files with datecode 001109 in the 001113 directory and/but they are SSD data with no counts so ignoring them. (I put these and the zero-length files into 001113/NA_files/ ) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL If data are available on tape reload the tape data into --> 001113t <-- directory which exists and is empty now. - Completed analysis with the few missing files. 001114 [2085]1-156,171-187: - HD2341[D] and HF2517[C] (Brookhaven 3/96; damaged at alignement) - A quick look at the SSD data from the non-zero 001113 files shows SSD is working. The gain was a little (~10%) low so I checked the GAIN setting on the 6300 amplifier: it was set at 9.10 whereas back in May '99 we set it to 10.00 -- **--> I set it to 9.90 so that tonight's SSD data (taken near the end) will be better calibrated. :( Zero-length or missing PSPC items 158-160, SMD 157-160; SSD items 171-188 found in 001115 and moved to 001114. Put zero-length files in '/NA_files/ - Analysis: SUMMARY, SMD, SSD, PSPC_CAL If data are available on tape reload the tape data into --> 001114t <-- directory which exists and is empty now. - Completed analysis with the few missing files. 001115 [2087]1-187: - HE2412[E] and HE2426[F] (---; ---) - Looks OK !?! - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT Grating/ Test 1 keV 4 keV 8 keV HE2412/001115 p0: 5.58 26.01 14.33 HE2412/951124 p0: 5.44 24.98 13.78 /nfs/bodicea/d4/F006H/951124 HE2426/001115 p0: 3.61 18.36 9.56 HE2426/960516 ave: 3.89 17.79 9.06 001116 [2089]1-191: - MA1022[A] and MA1026[A] (---; ---) - First MEG11 test run. - Looks OK !?! - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT Grating/ Test 1 keV 4 keV 8 keV MA1022/001116 p0: 12.46 23.48 10.19 MA1022/950628 p0: 11.60 23.39 10.17 /nfs/bodicea/d2/F001M/950628 MA1026/001116 p0: 12.08 24.19 10.68 MA1026/950624 p0: 11.69 23.60 10.45 /nfs/bodicea/d2/F001M/950624 001117 [2091]1-191: - MB1148[D] and MB1158[D] (Brookhaven 3/96; ---) - Looks OK !?! - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT Grating/ Test 1 keV 4 keV 8 keV MB1148/001117 p0: 17.08 20.90 8.27 MB1148/960122 p0: 14.70 22.14 9.16 /nfs/bodicea/d4/F009M/960122 MB1158/001117 p0: 14.75 22.92 9.57 MB1158/960126 p0: 12.59 25.06 11.10 /nfs/bodicea/d2/F009Mb/960126 :( 001121: [2100]1-191: - MC1200[C] and ME1409[E] (---; ---) - Analysis: SUMMARY, SMD, SSD :( Oh man, the PSPC Y voltages go from -0.6 to -0.4 V instead of -0.250 to +5.0 -- looks like something happened to the Y-axis circuitry... The X-axis looks OK with diffracted peaks visible... - Use the SSD measurements with 001127: 001127: [2103]1-136, [2100]137-190: - MC1200[C] and ME1409[E] (---; ---); - So far, so good... Ooops: SSD data has zero counts... !- Copy the SSD tests from 001121 into 001127 to have a complete test of these gratings... - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks Good! Grating/ Test 1 keV 4 keV 8 keV MC1200/001127 p0: 16.92 19.59 7.66 MC1200/951002 p0: 15.11 20.56 8.38 /nfs/bodicea/d3/F005M/951002 ME1409/001127 p0: 20.35 21.49 8.23 ME1409/951222 p0: 17.28 23.50 9.52 /nfs/bodicea/d2/F007M/951222 001207: [2113]1-190: - MA1012[B] and MA1025[A] (---; ---); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks Good! Grating/ Test 1 keV 4 keV 8 keV MA1012/001207 p0: 11.88 22.37 9.70 MA1012/950610 p0: 11.54 22.60 9.87 /nfs/bodicea/d2/F001M/950610 MA1025/001207 p0: 11.09 25.16 11.54 MA1025/950628 p0: 11.06 25.24 11.45 /nfs/bodicea/d2/F001M/950628 001211: [2117]1-190: - MB1165[D] and MD1367[F] (VSG gratings); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks OK. Grating/ Test 1 keV 4 keV 8 keV MB1165/001211 p0: 15.03 22.56 9.33 MB1165/990616 p0: 15.16 22.52 9.30 /nfs/spectra/d1/990616 MD1367/001211 p0: 15.34 21.52 8.85 MD1367/990623 p0: 15.97 22.01 9.11 /nfs/spectra/d1/990623 001212: [2119]1-188: - HC2251[C] and HD2334[D] (VSG gratings); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks OK. Grating/ Test 1 keV 4 keV 8 keV HC2251/001212 p0: 6.36 24.69 12.99 HC2251/990719 p0: 6.34 24.73 12.96 /nfs/spectra/d1/990719 HD2334/001212 p0: 5.40 26.53 15.27 HD2334/990719 p0: 5.70 26.23 15.02 /nfs/spectra/d1/990719 001213: [2121]1-187: - HA2003[A] and HA2045[A] (VSG gratings); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks OK. Grating/ Test 1 keV 4 keV 8 keV HA2003/001213 p0: 7.96 30.84 16.58 HA2003/990713 p0: 7.68 30.67 16.50 /nfs/spectra/d1/990713 HA2045/001213 p0: 6.69 23.79 11.84 HA2045/990713 p0: 6.87 23.63 11.74 /nfs/spectra/d1/990713 001214: [2123]1-188: - HF2503[] and HE2423[] (VSG gratings); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks OK. Grating/ Test 1 keV 4 keV 8 keV HF2503/001214 p0: 8.12 26.08 12.81 HF2503/990721 p0: 7.76 26.61 13.13 /nfs/spectra/d1/990721 HE2423/001214 p0: 4.12 15.33 7.35 HE2423/990721 p0: 3.67 15.48 7.61 /nfs/spectra/d1/990721 001218: [2125]1-191: - MA1019[] and MA1040[] (VSG gratings); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks OK. Grating/ Test 1 keV 4 keV 8 keV MA1019/001218 p0: 10.85 20.73 9.02 MA1019/990622 p0: 10.68 21.03 9.21 /nfs/spectra/d1/990622 MA1040/001218 p0: 12.48 26.74 12.30 MA1040/990622 p0: 12.38 26.80 12.37 /nfs/spectra/d1/990622 001221: [2127]1-191: - MC1212[] and ME1408[] (VSG gratings); - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Looks OK. - The files have '1409 instead of '1408 - modify the X-GEF analysis routine xgef_grat_name for the change. Grating/ Test 1 keV 4 keV 8 keV MC1212/001221 p0: 18.04 20.54 8.00 MC1212/990616 p0: 17.87 20.94 8.21 /nfs/spectra/d1/990616 ME1408/001221 p0: 18.57 21.60 8.46 ME1408/990623 p0: 18.46 22.06 8.68 /nfs/spectra/d1/990623 [ME1409/951222 p0: 17.28 23.50 9.52 /nfs/bodicea/d2/F007M/951222 ] ============================================= - - - RADIATION EXPOSURES HAPPENED HERE - - - ============================================= 010301: [2145]1-191: - MA1025[] and MA1026[] (RadEx,-; RadEX,-) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Analysis Looks OK. Grating/ Test 1 keV 4 keV 8 keV BarFrac/RectThick AmpF MA1025/010301 p0: 11.69 25.24 11.40 0.61/4213 1.044 MA1025/001207 p0: 11.09 25.16 11.54 0.61/4272 1.030 MA1025/950628 p0: 11.06 25.24 11.45 /nfs/bodicea/d2/F001M/950628 MA1026/010301 p0: 12.15 24.34 10.75 0.62/4066 1.060 MA1026/001116 p0: 12.08 24.19 10.68 0.62/4081 1.057 MA1026/950624 p0: 11.69 23.60 10.45 /nfs/bodicea/d2/F001M/950624 010302: [2147]1-191: - MB1148[] and MB1158[] (RadEx,Brookhaven 3/96; RadEX,-) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Analysis Looks OK. Grating/ Test 1 keV 4 keV 8 keV BarFrac/RectThick AmpF MB1148/010302 p0: 17.29 20.52 8.07 0.49/3353 1.088 MB1148/001117 p0: 17.08 20.90 8.27 0.49/3405 1.095 MB1148/960122 p0: 14.70 22.14 9.16 /nfs/bodicea/d4/F009M/960122 MB1158/010302 p0: 14.87 22.77 9.53 0.47/3901 1.105 MB1158/001117 p0: 14.75 22.92 9.57 0.47/3848 1.068 MB1158/960126 p0: 12.59 25.06 11.10 /nfs/bodicea/d2/F009Mb/960126 010307: [2149]1-191: - MA1022[] and ME1409[] (RadEx,-; RadEX,-) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Analysis Looks OK. Grating/ Test 1 keV 4 keV 8 keV BarFrac/RectThick AmpF MA1022/010307 p0: 12.49 23.28 10.07 0.60/3929 1.021 MA1022/001116 p0: 12.46 23.48 10.19 0.60/3957 1.034 MA1022/950628 p0: 11.60 23.39 10.17 /nfs/bodicea/d2/F001M/950628 ME1409/010307 p0: 19.83 21.48 8.27 0.48/3286 1.124 ME1409/001127 p0: 20.35 21.49 8.23 0.48/3248 1.141 ME1409/951222 p0: 17.28 23.50 9.52 /nfs/bodicea/d2/F007M/951222 010308: [2151]1-188: - HC2214[] and HC2259[] (RadEx, damaged at alignment; RadEX, Eng HESS) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Analysis Looks OK. Grating/ Test 1 keV 4 keV 8 keV BarFrac/RectThick AmpF HC2214/010308 p0: 7.72 23.22 10.93 0.56/4477 0.949 HC2214/001101 p0: * 7.86 23.56 11.15 0.57/4542 0.943 HC2214/960207 ave: 7.81 23.42 11.08 0.55/4558 -.--- (not on disk) *missing a few PSPC files, no SSD HC2259/010308 p0: 5.33 27.15 15.51 0.55/5531 0.961 HC2259/001101 p0: * 5.79 27.29 15.46 0.55/5527 0.945 HC2259/960307 ave: 5.46 26.17 14.63 0.56/5516 -.--- (not on disk) *missing a few PSPC files, no SSD 010309: [2153]1-188: - HB2104[] and HE2426[] (RadEx, Synch. tested; RadEX,-) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Analysis Looks OK. Grating/ Test 1 keV 4 keV 8 keV BarFrac/RectThick AmpF HB2104/010309 p0: 6.66 30.20 17.56 0.55/5725 0.978 HB2104/001113 p0: 6.58 30.61 17.96 0.56/5782 1.000 HB2104/960730 ave: 6.14 29.65 17.40 0.57/5714 -.--- (not on disk) HE2426/010309 p0: 4.11 18.51 9.52 0.59/4974 0.802 HE2426/001115 p0: 3.61 18.36 9.56 0.62/4774 0.968 HE2426/960516 ave: 3.89 17.79 9.06 0.61/4706 -.--- (not on disk) 010312: [2155]1-188: - HE2412[] and HF2517[] (RadEx,-; RadEX, damaged at alignement) - Analysis: SUMMARY, SMD, SSD, PSPC_CAL, PSPC_ANAL, PSPC_FITS, AMP, TILT - Analysis Looks OK. Grating/ Test 1 keV 4 keV 8 keV BarFrac/RectThick AmpF HE2412/010312 p0: 5.82 25.93 14.24 0.57/5331 0.934 HE2412/001115 p0: 5.58 26.01 14.33 0.57/5346 0.949 HE2412/951124 p0: 5.44 24.98 13.78 /nfs/bodicea/d4/F006H/951124 HF2517/010312 p0: 8.33 25.20 12.00 0.57/4639 0.979 HF2517/001114 p0: 8.21 25.51 12.26 0.57/4654 1.020 HF2517/960428 ave: 7.95 24.80 12.00 0.55/4731 -.--- (not on disk) ====================================================================== CHRONOLOGICAL NOTES ------------------- 11/20 Started this file of status/notes. Run xgef_analysis,/SUMMARY on all the directories and move files as needed (leaving backup directories alone.) Look at tilt test data, compare test item #69 for HB2121 in 001030[2077] run and 990715[1992] runs: the bright order is on plus instead of minus side of PSPC X. Looks like Y is flipped also as HB2121 and HX220 (#33, MX078 is #34) illuminates the "top" five wires instead of the "bottom" (lower Y voltage) five wires in previous tests. 11/21 001120 run atempt: :( Mac memory error around test item #6 Analyze 001030 data with PSPC calibration flipped in sign. Had to fix pspc_xcal_shift, pspc_show_cal[1] to take into account the V to X slope... 001030 PSPC cal and analysis looks OK! (HEG7 test.) "Qualify" the system based on this and: Analyze and "qualify" 001117 data set also, MEG11 test. Process all data sets as far as possible. 11/22 All PSPC_CAL's through 001117 look OK. 001121 data set is in and has PSPC Y axis problem, "squashed" in Y direction to -0.6 to -0.4 V range instead of -0.25 - 5.0V range. arg! 11/27 PSPC data look OK for tonight's run. 11/28 Analyze the 001127 data set... SSD had failed during the test, BUT can use the 001121 SSD measurements! OK. 12/08 New data set: 001207. Start analysis... 12/11 Check 001207 analysis results. Enter info from last previous testing, FYI. Some of these had fewer lines, or no SSD, etc. 12/12 Analyze 001211 data... 12/13 Analyze 001212 data... 12/14 Post 001212 summary results. Analyze 001213 data... 12/15 Post 001213 summary results. Analyze 001214 data... 12/19 001214 looks OK; start 001218 anlaysis... 1/2/01 001218 is OK. 001221 has strange '1409p0 fit to the (OK-looking) data points? Turns out the Ti-K 2nd order had NaN for its value - add this criteria to the jfit_lnf_valid routine and redo /PSPC_FITS,/AMP. (Also, John moved his s/w so point to the new location in jfit_analysis.pro.) Looks like '1409 is really '1408 (which is the VSG grating)... the data seem to imply this as well... 1/3/01 Modify xgef_grat_name routine for the ME1409--> ME1408 name switch and re-analyze. Now MC1212 has anomalous large second order Ti effic.s; catch this by modifing jfit_lnf_valid to ignore effic.s above 10.0 also. 1/23/01 Copy the 7 VSG data sets to /nfs/spectra/d1 with the other VSG data. Before: /dev/dsk/c1t0d0s7 17711418 10073253 7461051 58% /spectra/d1 After: /dev/dsk/c1t0d0s7 17711418 12097440 5436864 69% /spectra/d1 about 290 Mb per test - OK. 2/27/01 From Steve O'Dell at ACIS telecon: - HETG facets. exposed to 10 nano Coul/sq cm 40-640 keV. Now they go to MIT for characterization. 3/5/01 Start analyzing new X-GEF tests 010301, 010302 (initially in files '02 and '05, moved to correct ones). 3/7/01 The 010301 and 010302 tests analyzed OK. Send out rad_note_010307.txt 3/9/01 Analyze 010307,'08... 3/12/01 Check cal of '07,'08 tests... Check '09's cal. Finish analysis of these three... Get "eff" from my bin/sun4 . 010308 HC2259 analysis shows no zeroth-order plot in jfit_compare.ps ??? - modify jfit_compare to catch NANs! 3/13/01 Analyze 010312 data... Reanalyze '07-'13 with new jfit_compare. Send out rad_note_010313.txt . 3/20/01 Meet with Mike, decide to look at individual effic. values before/after Work on s/w to make before/after plots, rad_loop.pro calling rad_compare.pro. needed to modify fits/lnf_mk_effics.pro so that pspc_effics.lnf had no missing lines, instead bad values are flagged with negative error. Reprocess (xgef_analysis,/PSPC_FITS,/AMP) tests as needed... v010302, v001117, v001127, v001115 3/21/01 Continue with rad_compare: adjust effic.lnf errors to remove the ref grating effic error; put prelim plots for 8 gratings in Rad_anal dir. 3/22/01 finish first pass on rad_compare and rad_loop; re-arrange rad files; create local rad web page; 3/26/01 Analyze most recent pairs of VSG tests using rad_compare for comparison with rad analyses... Need to re-analyze some to have all lines in pspc_effics.lnf: 001221 MC1212 post 001221 ME1408 post 001211 MB1165 post 001030, 990715 HB2121 post, pre 001214 HE2423 post 001030, 990715 H2102 post, pre unique ones: 001221, 001211, 001030, 990715, 001214 3/27/01 Work on analyzing the 3 test dates that we need but have incomplete test data: 001101, 001113, 001114 001101: HEG7 missing 156-160, PSPC_CAL OK 001113: HEG7 missing 158-160, PSPC_CAL OK 001114: HEG7 missing 157-160, PSPC_CAL OK modified xrcf/pspc_analyze_fits.pro and '/pspc_read_events.pro to handle non-existant files gracefully... 3/28/01 OK, have results from all the rad exposed gratings now - time to look plots/values over and think... 4/2/01 Looking into rad result anomalies, starting with MA1025 zo. Looks like 001207 analysis had messed up Cu SSD fitting - re-do SSD fits and re-do /PSPC_FITS,/AMP on this directory. Also discover that SSD seems to be working with larger FWHM in the range 001207 to 001221; check all SSD fit Chi2 values. This was the time after the SSD failed during the 001127 test... Another bizarre spectrum: SSD MgO Direct for 010312 has a mystery bump between the O and Mg peaks - otherwise it's OK (Mg zo measurement is probably OK.) SSD at Ti 4.5 keV has large Chi2 values and NaN values... To be continued... 4/4/01 Other SSD large-Chi-squared values: 4.5/4.9 keV fits for 001116, 001117, 001127, 010301, 010302 tests - looks like SSD gain 3% low is messing up fitting(?!) -- but these values are not used in current analysis so put off fixing this for now. 1.254/0.525 keV fits for: 010309, two regions of HE2426 missing the measurement because SSD messes up, similar to 010312; See rad_ssd_010309_MgO.ps . Looking at MB1148p4, second order Ti-K: the second order is so low that the fits are meaningless - eliminate these measurements by requiring the FWHM < E/4 for a valid measurement. Fixes MB1158m8 gross 2nd order problem Check into MB1158p8 second order (Chi2=29.6)... Require FWHM > 0 ! Residual 2nd order Chi2=6.99 in p8 is due to poor fitting of Al-K data - actual pha spectra show good ROI counts agreement before/after. Assemble list of "anomalies" in the rad_compare plots... 5/1/01 Continue anomaly investigations... Several gratings show 0.93 keV first-order anomalies: MB1148, HC2259, HF2517. Looking at the 'anal_plot.ps files the ratio of Cu first-order fitted counts to ROI counts is generally around 0.8 +/-0.02; but for some the ratio is ~ 0.69. This fitting error (due to Cu line being double?!?) contributes to plus/minus asymetry and causes it to be consistant across regions if the error happens with the ref grating fitting, as it does in HF2517/001114 and HC2259/001101 measurements. Looking at MB1148... 5/2/01 - - - MB1148, Cu first order: measure fit-counts/ROI-counts for MEG Cu line... It's in the range 0.77-0.83 which seems OK and consistant enough. Because all the MB1148/001117 regions are low (and MB1158 ones too) it is likely that the Ref "pre, plus" measurement is High. However the Ref minus/plus ratio for 001117 seems in range for MEG tests: 001116 1.13 001117 1.15 001127 1.17 001207 1.16 +/- 2.2% stat error 010301 1.23 010302 1.25 010307 1.27 ... so it's hard to blame the reference for more than few per cent. See changes in Det(room) pressure and Src_Vac right around the MX078 Ref measurement ([2091]#129) - could PSPC internal gas have varied? Conclude that it's probably either: i) some ref/PSPC thing common to '1148 and '1158 OR ii) it may be a rad effect that BOTH '1148 and '1158 experienced. - - - For HEG gratings with Cu L (0.93 keV) problems: modify xrtf/pspc_analyze_fits.pro to include a ghost peak in the Cu L fit - looks a little better. Re-analyze 010312 and 001114 to see if this improves the comparison plots... 5/4/01 Re-analyzed all Rad HEG tests: 001101, 001113, 001114, 001115, 010308, 010309, 010312. The HEG Cu points look a little better... It may be that there are other systematic effects with the PSPC at Cu-L... 5/7/01 Make a table of the HEG plus/minus ratio at Cu-L to see if this varies in any strange way (e.g., can blame the reference measurement...) The HX220 Cu-L minus/plus counts ratio is in the range 1.03 to 1.17: 001101 1.17 001113 1.03 001114 1.15 001115 1.06 +/- 5.0% stat error 010308 1.14 010309 1.10 010312 1.11 For the pairs of tests we're using the ref difference is 7% at most... The variation seen could be due to statistics only?! Small bug found in rad_compare fixed - small effect on only one plot. Update the rad_anom.txt file (rad_anom.010507.txt) and send e-mail note on progress. 5/8/01 Start looking into 2nd-order anomalies... For MA1025 probably just a reference measurement stat error. For ME1409p8 and HC2259m8 there are clear mis-fits. 5/21/01 Look at HF2517 zeroth-order 5% effic change... Measurements/analysis look OK but from 4/2 above: "Another bizarre spectrum: SSD MgO Direct for 010312 has a mystery bump between the O and Mg peaks - otherwise it's OK (Mg zo measurement is probably OK.)" This might create a 5% effect in all regions... This effect was also seen in 010309 data which is the post-exposure measurement for HB2104, so the zeroth-order in '2104 could be this as well. HB2104 p4 region being 5% low at all first-order energies must be something in the sample as the grating/regions are scanned for each energy - would be hard to have an instruemental effect occur only for each time the p4 region was in the beam and not for other regions... The p8,m8 HB2104 regions seem to have more first-order variation as well... Best to look at the grating itself to a localized cause... Send around anomalies sheet. - - -